a car avrages 28.7 miles per gallon. how many gallons will it use on a trip 501.5 mmiles. round to the nearest tenth
Found 2 solutions by mananth, contestant:
Answer by mananth(16946)

70% male
30% female
...
7/3= x/24
x=24*7/3
x= 56 men
...
m.ananth@hotmail.ca

Answer by contestant(1)